A CD starts from rest at an angular acceleration of 3 radians/ s2.
(a) Find the angular speed after 10 seconds.
(b) Through what angle does the CD turn in this time of 10 seconds?
Answer #1

(a)

Angular acceleration (\alpha) = 3 rad / s^{2}

Initial angular velocity( \omega_{0} ​​​​​​) = 0

\omega (t) = \omega_{0} + \alpha t

\omega (t=10 sec) = 3 x 10 rad/sec = 30 rad/sec

This angular speed after 10 seconds is 30 rad/sec

(b)

\theta = \omega_{0} t + \alpha t^{2} /2

= 0 +( 3 x 10 x (10/2)) radians (\omega_{0} = 0)

= 150 radians

Thus CD turns by 150 radians